# Weighted algebraic topology, II

### 05 May 2026 v1

@@ Abstract.
Extending Lawvere’s ‘metric spaces’, we introduce relative metrics, with values in the extended real line. Formally, this ordered set is a symmetric monoidal closed category, and our structures are enriched categories on the latter. Concretely, the present goal is measuring ‘profits’ and ‘losses’ of a process, in any sense – possibly related to energy, or a variable in any science.
In particular, linear relative metrics derive from a potential function. Moreover, the spacetime of special relativity can be given a relative metric (or better a relative ‘antimetric’) derived from the metric tensor, much in the same way as a Riemannian manifold gets on ordinary metric.
@@
Generalised metric space, directed algebraic topology, weighted algebraic topology, Minkowski spacetime
@@ 0.1. An elementary example
(a) This article is Part II in a series which began with [G3], in 2007. The matter of the latter, reorganised in the general context of Directed Algebraic Topology, can also be found in Chapter 6 of the book [G4]. Both texts are freely available.
A prime structure analysed in [G3] consists of Lawvere’s generalised metrics [Lw], with values in the extended positive line $\overline{\mathbb{R}}^{+}=[0,\infty]$. Here we further extend this notion, introducing ‘relative metrics’ with values in the extended line $\overline{\mathbb{R}}=[-\infty,\infty]$, to express ‘profits’ and ‘losses’, in any sense – possibly related to a variable of any science, from Physics to Economy.
@@
(b) We consider the vertical profile $a$ of a track, or road (in a limited area of the Earth), represented in the $xy$-plane, with points $p^{\prime}=(x^{\prime},y^{\prime})$, $p^{\prime\prime}=(x^{\prime\prime},y^{\prime\prime})$, $p_{j}=(x_{j},y_{j})$
@@
As usual, its horizontal projection on the $x$-axis represents the ‘rectified horizontal path’, that is, the rectification of its projection on a horizontal plane, where the distance is measured by $\delta_{1}(p_{1},p_{2})=|x_{2}-x_{1}|$. The vertical axis measures the altitude, with distance $\delta_{2}(p,q)$ = $|y_{2}-y_{1}|$.
@@
(c) We are also interested in distinguishing ascent and descent – supposing we are going from $p^{\prime}$ to $p^{\prime\prime}$. This is not measured by the pseudometric $\delta_{2}$, but can be measured by a ‘relative metric’ $\mu$, with real values, and its positive part $\mu^{+}=\mu{\,{\scriptstyle{\vee}}\,}0$
(3) $\mu(p_{1},p_{2})=y_{2}-y_{1},\qquad\;\;\mu^{+}(p_{1},p_{2})=\max(y_{2}-y_{1},0)\geqslant 0.$
Let us note that $\mu$ is ‘linear’ (it is finite and satisfies a triangular identity), and expresses the increment of the gravitational potential per mass-unit, while $\mu^{+}$ is a quasi-pseudo metric [Ke] and a generalised metric in Lawvere’s sense [Lw] (see 0.2). Both $\mu$ and $\mu^{+}$ are invariant up to translation and not symmetric.
(d) The relative metric $\mu$ gives a relative valuation $v_{\mu}(a)$ of the path, which – because of linearity – is simply the increment in altitude from $p^{\prime}$ to $p^{\prime\prime}$
(4) $v_{\mu}(a)={\rm sup}_{\underline{p}}\,(\raisebox{1.93747pt}{${\mbox{$\sum$}}$}_{j}\;\mu(p_{j-1},p_{j}))=y^{\prime\prime}-y^{\prime}\in\mathbb{R},$
@@
The total ascent $v^{+}_{\mu}(a)$ of the path is defined as its $\mu^{+}$-valuation
(5) $\begin{array}[]{c}v^{+}_{\mu}(a)=v_{\mu^{+}}(a)={\rm sup}_{p}\,(\raisebox{1.93747pt}{${\mbox{$\sum$}}$}_{j}\;\mu^{+}(p_{j-1},p_{j})),\\[5.0pt] v^{+}_{\mu}(a)\,\in\,\mathbb{R}^{+}=[0,\infty[,\qquad v^{+}_{\mu}(a)\geqslant v_{\mu}(a).\end{array}$
We assume that the track is the graph of a piecewise monotone function $a\colon[x^{\prime},x^{\prime\prime}]\to\mathbb{R}$. Thus $v^{+}_{\mu}(a)$ is finite: the sum of increments in altitude on the maximal intervals on which the function $a$ is increasing.
The total descent $v^{-}_{\mu}(a)$, defined as
(6) $v^{-}_{\mu}(a)=v^{+}_{\mu}(a)-v_{\mu}(a)\geqslant 0\qquad\;\;(v_{\mu}(a)=v^{+}_{\mu}(a)-v^{-}_{\mu}(a)),$
is also a real number. Owing again to the linearity of $\mu$, $v^{-}_{\mu}(a)$ can also be obtained as the total ascent of the reversed path $a^{\sharp}$, going backwards (as proved in Lemma 5.3)
(7) $v^{+}_{\mu}(a^{\sharp})={\rm sup}_{p}\,(\raisebox{1.93747pt}{${\mbox{$\sum$}}$}_{j}\;\mu^{+}(p_{j},p_{j-1})).$
The absolute value $|\mu|(p_{1},p_{2})=|y_{2}-x_{2}|$ is here a pseudometric (another consequence of the linearity of $\mu$). The $|\mu|$-valuation of the path expresses the total variation of altitude
(8) $v_{|\mu|}(a)={\rm sup}_{p}(\raisebox{1.93747pt}{${\mbox{$\sum$}}$}_{j}|\mu|(p_{j-1},p_{j})),\qquad v_{|\mu|}(a)=v^{+}_{\mu}(a)+v^{-}_{\mu}(a).$
(e) Figure (1) might also represent the value $y$ of any variable (in Physics, Chemistry, Natural Sciences, Economy, Sociology, etc.) as a function of the variable $x$ (position on a line, or time, etc.). The interpretation of $v_{\mu}(a)$, $v^{+}_{\mu}(a)$, $v^{-}_{\mu}(a)$, $v_{|\mu|}(a)$ is similar.
@@ 0.2. Positive metrics and weighted algebraic topology
Lawvere’s article [Lw] presented, in 1974, a (generalised) metric space $X$ as a small category enriched on the positive extended real line ${\bf w}^{+}=\overline{\mathbb{R}}^{+}=[0,\infty]$; the latter is viewed as a category (with morphisms $\lambda\geqslant\mu$) endowed with a strict symmetric monoidal closed structure given by the (extended) sum $\lambda+\mu$.
@@
Similar extensions have been studied in the theory of metric spaces, like quasi-pseudo metrics [Ke], which are finite, or écarts in Bourbaki [Bk], which are symmetric.
Lawvere’s extension combines a formal elegance within category theory (one might say ‘necessity’) with concrete advantages, e.g. with respect to the existence of limits and colimits – which obviously fails for ‘finite metrics’. (An influential part of the article [Lw], developed by several authors, extends part of the theory of metric spaces, like Cauchy completeness, to enriched categories.)
Breaking the symmetry property, such a metric can also express privileged directions. It was used in this sense by the present author, under the name of $\delta$-metric, as a base for an enriched version of Directed Algebraic Topology [G3, G4]. Lipschitz paths and Lipschitz homotopies are based on the standard $\delta$-interval $\delta\mathbb{I}=\delta[0,1]$, with $\delta$-metric
@@
The length of a Lipschitz path $a\colon\delta\mathbb{I}\to X$ is given by the following least upper bound in $\overline{\mathbb{R}}^{+}$
@@
A $\delta$-metric space $X$ has a fundamental weighted category ${\rm w}\Pi_{1}(X)$ [G3, G4], an enrichment of the fundamental category of a directed spaces [G1, G4]: it has for objects the points of $X$, and for arrows $\xi\colon x\to x^{\prime}$ the classes of Lipschitz paths $a\colon\delta\mathbb{I}\to X$ from $x$ to $x^{\prime}$, up to the equivalence relation generated by Lipschitz homotopies with fixed endpoints. The weight of such an arrow $\xi\colon x\to x^{\prime}$ is defined as usual in a quotient, as the greatest lower bound of the length of the paths belonging to this equivalence class
@@
This gives ${\rm w}\Pi_{1}(X)$ a weight, or generalised norm, in $\bf w$$^{+}$ [Lw, G2, G3, G4]. All this is extended to the more general framework of ‘spaces with weighted paths’, whose finer quotients can interpret structures of Noncommutative Geometry: see [G3], Section 7, and [G4], Section 6.7.
0.3. Relative metrics
We now replace $\bf w$$^{+}$ by a larger structure $\bf w$} based on the extended real line $\overline{\mathbb{R}}=[-\infty,\infty]$. This complete lattice is again viewed as a category, with morphisms $\lambda\geqslant\mu$, and endowed with a strict symmetric monoidal closed structure given by the extended sum $\lambda+\mu$; here we let $\infty+\infty=\infty$ (adding any element must preserve the initial object $\infty$).
A $\mu$-metric space, or relative metric space, is thus a small category $X$ enriched on $\bf w$}. This means a set $X$ with a mapping $\mu\colon X\!\times\!X\to\overline{\mathbb{R}}$ that satisfies the same conditions as $\delta$ in (9)
(13) $\mu(x,y)+\mu(y,z)\geqslant\mu(x,z),\qquad\;\;0\geqslant\mu(x,x)\qquad(x,y,z\in X),$
but the second condition amounts here to $\mu(x,x)=0$ or -$\infty$, for all $x\in X$ (see 2.1).
The extended real number $\mu(x,y)$ expresses a relative measure of the transition from $x$ to $y$. The ‘valuation’, or ‘relative length’, of a path is also an extended real number (in 5.2).
The basic example is now the standard $\mu$-line $\mu\mathbb{R}$, with $\mu$-metric
(14) $\mu(x,y)=y-x.$
More generally, a linear $\mu$-metric on a set $X$ (finite and satisfying the ‘triangular identity’) can be expressed as the variation $\mu(x,y)=\varphi(y)-\varphi(x)$ of an arbitrary function $\varphi\colon X\to\mathbb{R}$, called the potential of $\mu$ (see 3.2).
Linear relative metrics are important here (and trivial in the positive case); they could also be called ‘conservative metrics’, being related to a potential. These two features will be distinguished in Part III: the more general structure of a ‘space with valued paths’ presents a linear case (possibly determined by a 1-form on a smooth manifold, or a vector field on a Riemannian manifold), which only corresponds to a linear $\mu$-metric in the conservative case.
@@ 0.4. Outline
Section 2 introduces (relative) $\mu$-metric spaces, developing their links to the positive case of $\delta$-metric spaces previously studied in [Lw, G3, G4] and other works. Limits and colimits are dealt with in 2.3. The main properties, like linearity and invariance up to group actions, are outlined in 2.2 and studied in Section 3. As a critical problem, quotients of $\mu$-spaces tend to be trivial: the line $\mu\mathbb{R}$ induces a chaotic $\mu$-metric on the circle (with constant value -$\infty$), see 3.1(c).
Section 4 deals with the reflective and coreflective symmetrisations of a $\mu$-space, the associated topologies, and the associated reflective and coreflective preorders. In different situations it is convenient to investigate spaces and maps by ‘reflective’ or ‘coreflective tools’ – another critical aspect discussed below. (In Part I only the reflective procedures were needed, for $\delta$-spaces.)
Paths in a $\mu$-space and their (relative) valuation are analysed in Section 5, extending the length of paths in a $\delta$-space. The tensor product of $\mu$-spaces is studied in Section 6, together with various examples, including relative metrics associated to a conservative vector field.
The last Section 7 briefly introduces relative antimetrics, that satisfy the reverse triangular inequality, and shows how the Minkowski spacetime of special relativity can be endowed with such a structure, derived from the metric tensor.
@@ 0.5. Comments and sequels
Weighted Algebraic Topology, as proposed in [G3, G4], is an enrichment of this topic: the truth-valued approach of directed algebraic topology (where a path is either allowed or not) is replaced with a measure of costs. The general aim is, now, measuring the cost of (possibly non-reversible) phenomena.
@@
(b) The second point is the more intriguing: relative metrics do have a topological content, studied throughout this paper, but a complex relation to topology. We should not expect a relative metric space to have a definite topology: there is rather a range of them, from the reflective to the coreflective topology (see Section 4), and even this range can be insufficient, as in 4.8. The main interest of relative metrics seems to be to evaluate paths (in Section 5).
A convincing evidence in this sense is related to the Minkowski spacetime: the path of a material or lightlike particle is continuous for the natural (euclidean!) topology, which has little to do with the metric tensor, and the $\mu$-metric derived here from the latter (as discussed in 7.5 and 7.6).
(c) Part III will overcome these problems, and others, introducing the more flexible setting of valued spaces, or spaces with valued paths – a relative version of the positive case already studied in [G3, G4]. In this framework an object already bears a given topology, and quotients work as one can expect: for instance, the valued line $\rm v$$\mathbb{R}$ (based on the present $\mu$-line $\mu\mathbb{R}$) has an appropriate quotient up to integral translations: the (linearly) valued circle ${\rm v}\mathbb{S}^{1}$, where a path is valued by any lifting to the line, and a loop is valued by its (integral) winding number; this also shows that – in this context – the linear case need not be conservative, as already mentioned in 0.3.
(d) A sequel will develop homotopy theory for $\mu$-spaces and valued spaces, introducing and studying their ‘fundamental valued category’ (equipped with a relative valuation).

### 13 May 2026 v2

@@ Abstract.
Extending the ‘metric spaces’ of Lawvere, we study ‘real metrics’, with values in the extended real line $\infty$$\infty$].
Formally, this ordered set is a symmetric monoidal closed category, and our structures are enriched categories on the latter. Concretely, the present goal is measuring ‘profits’ and ‘losses’ of a process, in any sense – possibly related to energy, or a variable in any science. In particular, linear real metrics derive from a potential function.
@@
Generalised metric space, directed algebraic topology, enriched categories, weighted algebraic topology
@@ Introduction
0.1. Metrics with real values
This article is Part II of a series which began with [G3], in 2007, although at the time a sequel was not foreseen. The matter of the latter, reorganised in the general context of Directed Algebraic Topology, can also be found in Chapter 6 of the book [G4]. Both texts are freely available.
A prime structure analysed in [G3] consists of Lawvere’s generalised metrics [L1], with values in the extended positive line $\overline{\mathbb{R}}^{+}=[0,\infty]$. Here we further extend this notion, studying ‘real metrics’, with values in the extended real line $\overline{\mathbb{R}}=[-\infty,\infty]$, to express ‘profits’ and ‘losses’, in any sense – possibly related to a variable of any science, from Physics to Economy.
Two unpublished preprints of 1984 and 2015 also use metrics with values in $\overline{\mathbb{R}}$, in different perspectives. Dealing with entropy, Lawvere introduced a real valued antimetric $M(x,y)$ on the class of objects of a category $\mathbb{X}$ endowed with an ‘entropy supply’ $s\colon\mathbb{X}\to\mathbb{R}$ ([L2], p. 4). Dealing with the Legendre-Fenchel transform for real vector spaces, Willerton used real valued metrics, called $\overline{\mathbb{R}}$-metrics [Wi].
@@ 0.2. An elementary example
(a) We consider the vertical profile $a$ of a track, or road (in a limited area of the Earth), represented in the $xy$-plane, with points $p^{\prime}=(x^{\prime},y^{\prime})$, $p^{\prime\prime}=(x^{\prime\prime},y^{\prime\prime})$, $p_{j}=(x_{j},y_{j})$
@@
As usual in such maps (for mountain paths, etc.), the horizontal projection on the $x$-axis represents the ‘rectified horizontal path’, that is, the rectification of the projection of the track on a horizontal plane, where the distance is measured by $\delta_{1}(p_{1},p_{2})=|x_{2}-x_{1}|$. The vertical axis measures the altitude, with distance $\delta_{2}(p,q)$ = $|y_{2}-y_{1}|$.
@@
(b) We are also interested in distinguishing ascent and descent – supposing we are going from $p^{\prime}$ to $p^{\prime\prime}$. This is not measured by the pseudometric $\delta_{2}$, but can be measured by a a real valued metric $\rho$, and its positive part $\rho^{+}=\rho{\,{\scriptstyle{\vee}}\,}0$
(3) $\rho(p_{1},p_{2})=y_{2}-y_{1},\qquad\;\;\rho^{+}(p_{1},p_{2})=\max(y_{2}-y_{1},0)\geqslant 0.$
Let us note that $\rho$ is ‘linear’ (it is finite and ‘additive’, see 2.2(c)), and expresses the increment of the gravitational potential per mass-unit, while $\rho^{+}$ is a quasi-pseudo metric [Ke] and a generalised metric in Lawvere’s sense [L1] (see 0.3). Both $\rho$ and $\rho^{+}$ are invariant up to translation and not symmetric.
(c) The real metric $\rho$ gives a real valued valuation $v_{\rho}(a)$ of the path, which, by linearity, is simply the increment in altitude from $p^{\prime}$ to $p^{\prime\prime}$
(4) $v_{\rho}(a)={\rm sup}_{\underline{p}}\,(\raisebox{1.93747pt}{${\mbox{$\sum$}}$}_{j}\;\rho(p_{j-1},p_{j}))=y^{\prime\prime}-y^{\prime}\in\mathbb{R},$
@@
The total ascent $v^{+}_{\rho}(a)$ of the path is defined as its $\rho^{+}$-valuation
(5) $\begin{array}[]{c}v^{+}_{\rho}(a)=v_{\rho^{+}}(a)={\rm sup}_{p}\,(\raisebox{1.93747pt}{${\mbox{$\sum$}}$}_{j}\;\rho^{+}(p_{j-1},p_{j})),\\[5.0pt] v^{+}_{\rho}(a)\,\in\,\mathbb{R}^{+}=[0,\infty[,\qquad v^{+}_{\rho}(a)\geqslant v_{\rho}(a).\end{array}$
We assume that the track is the graph of a piecewise monotone continuous function $a\colon[x^{\prime},x^{\prime\prime}]\to\mathbb{R}$. Thus $v^{+}_{\rho}(a)$ is finite: the sum of increments in altitude on the maximal intervals on which the function $a$ is increasing.
The total descent $v^{-}_{\rho}(a)$, defined as
(6) $v^{-}_{\rho}(a)=v^{+}_{\rho}(a)-v_{\rho}(a)\geqslant 0\qquad\;\;(v_{\rho}(a)=v^{+}_{\rho}(a)-v^{-}_{\rho}(a)),$
is also a real number. Owing again to the linearity of $\rho$, $v^{-}_{\rho}(a)$ can also be obtained as the total ascent of the reversed path $a^{\sharp}$, going backwards (as proved in Lemma 5.3)
(7) $v^{+}_{\rho}(a^{\sharp})={\rm sup}_{p}\,(\raisebox{1.93747pt}{${\mbox{$\sum$}}$}_{j}\;\rho^{+}(p_{j},p_{j-1})).$
The absolute value $|\rho|(p_{1},p_{2})=|y_{2}-x_{2}|$ is here a pseudometric (another consequence of the linearity of $\rho$). The $|\rho|$-valuation of the path expresses the total variation of altitude
(8) $v_{|\rho|}(a)={\rm sup}_{p}(\raisebox{1.93747pt}{${\mbox{$\sum$}}$}_{j}|\rho|(p_{j-1},p_{j})),\qquad v_{|\rho|}(a)=v^{+}_{\rho}(a)+v^{-}_{\rho}(a).$
(d) Figure (1) might also represent the value $y$ of any variable (in Physics, Chemistry, Natural Sciences, Economy, Sociology, etc.) as a function of the variable $x$ (position on a line, or time, etc.). The interpretation of $v_{\rho}(a)$, $v^{+}_{\rho}(a)$, $v^{-}_{\rho}(a)$, $v_{|\rho|}(a)$ is similar.
@@ 0.3. Positive metrics and weighted algebraic topology
Lawvere’s article [L1] presented, in 1974, a (generalised) metric space $X$ as a small category enriched on the positive extended real line ${\bf w}^{+}=\overline{\mathbb{R}}^{+}=[0,\infty]$; the latter is viewed as a category (with morphisms $\lambda\geqslant\rho$) endowed with a strict symmetric monoidal closed structure given by the (extended) sum $\lambda+\rho$.
@@
Similar extensions have been studied in the theory of metric spaces, like the quasi-pseudo metrics of [Ke], which take finite values, or the écarts of Bourbaki [Bk], which are symmetric.
Lawvere’s extension combines a formal elegance within category theory (one might say ‘necessity’) with concrete advantages, e.g. with respect to the existence of limits and colimits – which obviously fails for ‘finite metrics’. (An influential part of the article [L1], developed by several authors, extends part of the theory of metric spaces, like Cauchy completeness, to enriched categories.)
Breaking the symmetry property, such a metric can also express privileged directions. It was used in this sense by the present author, under the name of $\delta$-metric, as a base for an enriched version of Directed Algebraic Topology [G3, G4]. Lipschitz paths and Lipschitz homotopies are based on the standard $\delta$-interval $\delta\mathbb{I}=\delta[0,1]$, with a $\delta$-metric meant to hinder any backward move
@@
The length of a Lipschitz path $a\colon\delta\mathbb{I}\to X$ is given by the following least upper bound in $\overline{\mathbb{R}}^{+}$:
@@
A $\delta$-metric space $X$, or $\delta$-space, has a fundamental weighted category ${\rm w}\Pi_{1}(X)$ [G3, G4], an enrichment of the fundamental category of a directed spaces [G1, G4]: it has for objects the points of $X$, and for arrows $\xi\colon x\to x^{\prime}$ the classes of Lipschitz paths $a\colon\delta\mathbb{I}\to X$ from $x$ to $x^{\prime}$, up to the equivalence relation generated by Lipschitz homotopies with fixed endpoints. The weight of such an arrow $\xi\colon x\to x^{\prime}$ is defined as usual in a quotient, as the greatest lower bound of the length of the paths belonging to this equivalence class
@@
This gives ${\rm w}\Pi_{1}(X)$ a weight, or generalised norm, in $\bf w$$^{+}$ [L1, G2, G3, G4]. All this is extended to the more general framework of ‘spaces with weighted paths’, whose finer quotients can interpret structures of Noncommutative Geometry: see [G3], Section 7, and [G4], Section 6.7.
0.4. Real valued metrics
We now replace $\bf w$$^{+}$ by a larger structure $\bf w$} based on the extended real line $\overline{\mathbb{R}}=[-\infty,\infty]$. This complete lattice is again viewed as a category, with morphisms $\lambda\geqslant\rho$, and endowed with a strict symmetric monoidal closed structure given by the extended sum $\lambda+\rho$; here we let $\infty+\infty=\infty$ (adding any element must preserve the initial object $\infty$).
A $\rho$-metric space, or $\rho$-space, is thus a small category $X$ enriched on $\bf w$}. This means a set $X$ with a mapping $\rho\colon X\!\times\!X\to\overline{\mathbb{R}}$, called a real metric, or $\rho$-metric, that satisfies the same conditions as $\delta$ in (9)
(13) $\rho(x,y)+\rho(y,z)\geqslant\rho(x,z),\qquad\;\;0\geqslant\rho(x,x)\qquad(x,y,z\in X),$
but the second condition amounts here to $\rho(x,x)=0$ or -$\infty$, for all $x\in X$ (see 2.1).
The extended real number $\rho(x,y)$ expresses a measure of the transition from $x$ to $y$. The ‘valuation’, or ‘real-valued length’, of a path is also an extended real number (see 5.2).
The basic example is now the standard $\rho$-line $\rho\mathbb{R}$, with real metric
(14) $\rho(x,y)=y-x.$
More generally, a linear $\rho$-metric on a set $X$ (see 2.2(c)) can be expressed as the variation $\rho(x,y)=\Phi(y)-\Phi(x)$ of an arbitrary function $\Phi\colon X\to\mathbb{R}$, called the potential of $\rho$ (see 3.2).
Linear real metrics are important here (and trivial in the positive case); they could also be called ‘conservative metrics’, being related to a potential. These two features will be distinguished in Part III: the more general structure of a ‘space with valued paths’ presents a linear case (possibly determined by a closed 1-form on a smooth manifold), which only corresponds to a linear $\rho$-metric in the conservative case (for an exact 1-form, or a conservative vector field on a Riemannian manifold).
@@ 0.5. Outline
Section 2 introduces $\rho$-metric spaces, developing their links to the positive case of $\delta$-metric spaces previously studied in [L1, G3, G4] and other works. Limits and colimits are dealt with in 2.3. The main properties, like linearity and invariance up to group actions, are outlined in 2.2 and studied in Section 3. As a critical problem, quotients of $\rho$-spaces tend to be trivial: the line $\rho\mathbb{R}$ induces a chaotic $\rho$-metric on the circle (with constant value -$\infty$), see 3.1(c).
Section 4 deals with the reflective and coreflective symmetrisations of a $\rho$-space, the associated topologies, and the associated reflective and coreflective preorders. In different situations it is convenient to investigate spaces and maps by ‘reflective’ or ‘coreflective tools’ – another critical aspect discussed below. (In Part I only the reflective procedures were needed, for $\delta$-spaces.)
Paths in a $\rho$-space and their valuation in $\overline{\mathbb{R}}$ are analysed in Section 5, extending the length of paths in a $\delta$-space. The tensor product of $\rho$-spaces is studied in Section 6, together with various examples, including real metrics associated to a conservative vector field.
@@ 0.6. Comments and sequels
Weighted Algebraic Topology, as proposed in [G3, G4], is an enrichment of this topic: the truth-valued approach of directed algebraic topology (where a path is either allowed or not) is replaced with a measure of costs, possibly infinite. The general aim is, now, measuring the cost of (possibly non-reversible) phenomena.
@@
(b) The second point is the more intriguing: real metrics do have a topological content, studied throughout this paper, but a complex relation to topology. We should not expect a real metric space to have a definite topology: there is rather a range of them, from the reflective to the coreflective topology (see Section 4), and even this range can be insufficient, as in 4.8. In our perspective, the main interest of real metrics is evaluating paths (see Section 5), but different goals have already appeared in [L2, Wi].
A convincing evidence in this sense will be explored in a sequel. The Minkowski spacetime of special relativity can be given a real valued metric, derived from the metric tensor, much in the same way as a Riemannian manifold gets an ordinary metric. The path of a material or lightlike particle is continuous for the natural (euclidean!) topology, which has little to do with the metric tensor, and the associated real metric. On the other hand, these structures do evaluate the proper duration of a causal path in the spacetime.
(c) Part III will overcome these problems, and others, introducing the more flexible setting of valued spaces, or spaces with valued paths – a real valued version of the positive case already studied in [G3, G4]. In this framework an object already bears a given topology, and quotients work as one can expect: for instance, the valued line $\rm v$$\mathbb{R}$ (based on the present $\rho$-line $\rho\mathbb{R}$) has an appropriate quotient up to integral translations: the (linearly) valued circle ${\rm v}\mathbb{S}^{1}$, where a path is valued by any lifting to the line $\rho\mathbb{R}$, and a loop is valued by its (integral) winding number; this also shows that, for valued spaces, the linear case need not be conservative, as already mentioned in 0.4.
Homotopy theory for $\rho$-spaces and valued spaces will also be studied.